# Operator-aware shadow importance sampling for accurate fidelity estimation

### v1

@@ Abstract
Estimating the fidelity between an unknown quantum state and a fixed target is a fundamental task in quantum information science. Direct fidelity estimation (DFE) enables this without full tomography by sampling observables according to a target-dependent distribution. However, existing approaches face notable trade-offs. Grouping-based DFE achieves strong accuracy for small systems but suffers from exponential scaling, and its applicability is restricted to Pauli measurements. In contrast, classical-shadow-based DFE offers scalability but yields lower accuracy on structured states. In this work, we address these limitations by developing two classes of operator-aware shadow importance sampling algorithms using informationally overcomplete positive operator-valued measures. Instantiated with local Pauli measurements, our algorithm improves upon the grouping-based algorithms for Haar-random states. For structured states such as the GHZ and W states, our algorithm also eliminates the exponential memory requirements of previous grouping-based methods. Numerical experiments confirm that our methods achieve state-of-the-art performance across Haar-random, GHZ, and W targets.
@@ II.2 Preliminaries on previous DFE protocols
that is, the normalized expectation value of the Pauli string $\mathbf{p}$. We denote the corresponding characteristic vector, indexed by Pauli strings, as $\bm{\chi}_{A}$. Suppose the unknown state is $\rho$, and let the target pure state have density matrix $O$. Then the fidelity can be expressed as
@@
G-DFE builds on this method by grouping qubit-wise commuting Pauli operators, thereby allowing the simultaneous estimation of multiple Pauli expectation values from a single measurement setting. Concretely, a measurement setting $\mathbf{p}$ is sampled from a subset of $\{I,X,Y,Z\}^{n}$ according to the distribution $p(\mathbf{p})$, where $\mathbf{p}$ represents a group of Pauli operators that (qubit-wise) commute with $\mathbf{p}$. Then $\mathcal{I}(\mathbf{p})$ is defined as the index set of the characteristic vector corresponding to the Pauli strings in the sampled group. In G-DFE, the probability of sampling $\mathbf{p}$ is given by $\left\|(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|^{2}$. Then the random variable associated with $\mathbf{p}$ is defined as
$R=\frac{(\bm{\chi}_{\rho})_{\mathcal{I}(\mathbf{p})}\cdot(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}}{\left\|(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|^{2}}=\frac{(\bm{\chi}_{\rho})_{\mathcal{I}(\mathbf{p})}\cdot(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}}{\tilde{p}(\mathbf{p})}.$ (3)
@@ III Optimization for general states
Input: POVM $\bm{\Pi}=\{\Pi_{U,\mathbf{b}}\}_{U,\mathbf{b}}$ and default distribution $p$
@@
$\begin{aligned} \text{minimize}\quad&\sum_{U}p(U)t_{U}\\ \text{subject to}\quad&-t_{U}\leq\omega_{U,\mathbf{b}}\leq t_{U},\quad\forall U,\mathbf{b},\\ &\sum_{U,\mathbf{b}}\omega_{U,\mathbf{b}}\Pi_{U,\mathbf{b}}=O.\end{aligned}$
@@
OASIS-GT provides a general solution for random targets. However, as will be demonstrated in Section V, applying OASIS-GT to highly structured states such as the GHZ and W states shows diminished performance when compared with previous algorithms such as G-DFE and C-DFE. This is because for these special states, the Pauli operators with nonzero probabilities can be grouped very efficiently, which eliminates the potential advantage of optimizations that do not rely on grouping (such as LP).
ALGORITHM 2: OASIS-ST for the GHZ state.
@@
     $U\leftarrow\bigotimes_{i=1}^{n}\mathbf{p}_{i}$; $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
@@
     $U\leftarrow\bigotimes_{i=1}^{n}\mathbf{p}_{i}$; $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
@@ IV Optimization for structured states
In this section, we present efficient and scalable algorithms for structured states, specifically the GHZ and W states. Heuristic algorithms such as G-DFE show scalability issues due to the exponential resource requirements in the grouping procedure and the memory needed to store the groups. To overcome these limitations, we propose more efficient estimators by leveraging a non-heuristic grouping strategy, collectively referred to as OASIS-ST (operator-aware shadow importance sampling for structured targets).
IV.1 OASIS-ST for the GHZ state
@@ V.1 Setting
We conducted numerical experiments on Haar-random, GHZ, and W states for systems of 3, 4, 5, and 6 qubits to demonstrate the performance of the proposed algorithms. For all targets, we evaluated the proposed OASIS-GT method, and for the GHZ and W states, we also applied OASIS-ST. In each experiment, the target density matrix is denoted by $O$, and the unknown state is modeled as a depolarized version
@@ V.2 MSE comparison
For OASIS-GT and C-DFE, the user can directly specify the number of measurement shots, and in all reported settings this number is matched to that of G-DFE (see Table 4 in Appendix F) for a fair comparison. In contrast, for OASIS-ST, the user specifies $\epsilon$ and $\delta$; the total number of shots then becomes a random variable, and is comparable on average to that of G-DFE. Since the MSE typically scales inversely with the number of shots, we apply a correction factor so that the reported MSEs for OASIS-ST reflect the same average number of shots as G-DFE.
Optimization for Haar-random states was performed using an IOC-POVM based on uniform Pauli measurements, ensuring a consistent basis with G-DFE, which also employs Pauli measurements. The results show that for Haar-random states, OASIS-GT consistently outperforms G-DFE in terms of MSE. For both the GHZ and W states, OASIS-ST achieves the best performance for systems of four qubits or larger. More importantly, the key advantage of OASIS-ST lies in its scalability. That is, it provides a grouping-based estimator that remains practical for large systems.
Table 1: Comparison of MSE (1e-4) for Haar-random states. C-DFE and OASIS-ST are inapplicable, as they are designed for structured targets.
@@
OASIS-ST 2.77 2.35 1.40 .721
V.3 Understanding the improvements in OASIS-ST
Fewer groups generally lead to smaller MSE, since more Pauli observables can be estimated from each measurement shot. In this sense, Algorithms 2 and 3 have the desirable property of minimizing the number of groups.
This grouping is valid, since all of these strings commute with $YYYZ$, but it is suboptimal and redundant because $YYYZ$ itself has zero probability. Consequently, this grouping does not minimize the number of groups, as shown in Table 5 in Appendix F. On the other hand, our OASIS-ST algorithm minimizes the number of groups while ensuring that each group contains no redundant Pauli strings, which accounts for the improvement achieved by our method.
@@ VI Discussion
We have introduced a framework for DFE based on operator-aware importance sampling. By formulating estimator optimization as a linear program over an IOC-POVM, OASIS-GT extends and improves upon existing DFE approaches. Furthermore, for structured targets such as the GHZ and W states, the proposed OASIS-ST provides improved fidelity estimation without exponential storage overhead.
A limitation of OASIS-GT is that, like G-DFE, it does not scale to many qubits for Haar-random states because most random states lack the structure necessary for an optimal DFE protocol to be formulated and implemented efficiently. Nevertheless, OASIS-GT provides a principled foundation for operator-aware importance sampling, and enhancing the scalability of the underlying optimization remains an important challenge.
While OASIS-ST demonstrates both strong performance and scalability, it still inherits a limitation in that it is currently applicable only to Pauli measurement settings. Extending the framework to accommodate more general measurement families remains an important direction for future work.
Additional promising directions include enhancing OASIS-GT through improved surrogate formulations and by considering other IOC-POVMs, and developing OASIS-ST beyond the GHZ and W states.
@@ Appendix B Derivation of Algorithm 1
Let $p(U)$ denote the default distribution of $U$ that defines $\bm{\Pi}$. In other words, we apply $\rho\mapsto U\rho U^{\dagger}$ with probability $p(U)$. Then the POVM elements are
@@
Let us denote the target state’s density matrix as $O$. Then our goal can be formulated as estimating $\text{tr}(\rho O)$ given many copies of an unknown state $\rho$. Since $\bm{\Pi}$ is an IOC-POVM, $O$ can be expressed as a linear combination of the elements of $\bm{\Pi}$ as
@@
minimize $\quad\sum_{U}p(U)\max_{\mathbf{b}}|\omega_{U,\mathbf{b}}|$ (5)
@@
minimize $\quad\sum_{U}p(U)t_{U}$ (6)
@@ Appendix C Derivation of Algorithm 2
$m_{\mathbf{p}}=\left\lceil\frac{2\left\|(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|_{1}^{2}}{\left\|(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|^{4}dl\epsilon^{2}}\ln\frac{2}{\delta}\right\rceil$ (11)
@@
In branch 2, $(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\in\mathbb{R}^{1}$ and $\left|(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right|=1/\sqrt{d}$, so
@@ Appendix E Derivation of Algorithm 3
In branch 2, $|\mathcal{I}(\mathbf{p})|=d/4$ and all entries of $(\bm{\chi}_{O})_{\mathcal{I}(\mathbf{p})}$ are $2/n\sqrt{d}$. Therefore,
@@ Appendix F Additional tables
The following tables provide additional numerical data mentioned in Sections V.2 and V.3.
@@
Table 5: Number of groups produced by G-DFE and OASIS-ST for the GHZ and W states.

### v2

@@ Abstract
Estimating the fidelity between an unknown quantum state and a fixed target is a fundamental task in quantum information science. Direct fidelity estimation (DFE) enables this without full tomography by sampling observables according to a target-dependent distribution. However, existing approaches face notable trade-offs. Grouping-based DFE achieves strong accuracy for small systems but suffers from exponential scaling, and its applicability is restricted to Pauli measurements. In contrast, classical-shadow-based DFE offers scalability but yields lower accuracy on structured states. In this work, we address these limitations by developing two classes of operator-aware shadow importance sampling algorithms using informationally overcomplete positive operator-valued measures. Instantiated with local Pauli measurements, our algorithm improves over the grouping-based algorithms for Haar-random states. For structured states such as the GHZ and W states, our algorithm also eliminates the exponential memory requirements of previous grouping-based methods. Numerical experiments confirm that our methods achieve state-of-the-art performance across Haar-random, GHZ, and W targets.
@@ II.2 Preliminaries on previous DFE protocols
that is, the normalized expectation value of the Pauli string $\mathbf{p}$. We denote the corresponding characteristic vector, indexed by Pauli strings, as $\boldsymbol{\chi}_{A}$. Suppose the unknown state is $\rho$, and let the target pure state have density matrix $O$. Then the fidelity can be expressed as
@@
G-DFE builds on this method by grouping qubit-wise commuting Pauli operators, thereby allowing the simultaneous estimation of multiple Pauli expectation values from a single measurement setting. Concretely, a measurement setting $\mathbf{p}$ is sampled from a subset of $\{I,X,Y,Z\}^{n}$ according to the distribution $p(\mathbf{p})$, where $\mathbf{p}$ represents a group of Pauli operators that (qubit-wise) commute with $\mathbf{p}$. Then $\mathcal{I}(\mathbf{p})$ is defined as the index set of the characteristic vector corresponding to the Pauli strings in the sampled group. In G-DFE, the probability of sampling $\mathbf{p}$ is given by $\left\|(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|^{2}$. Then the random variable associated with $\mathbf{p}$ is defined as
$R=\frac{(\boldsymbol{\chi}_{\rho})_{\mathcal{I}(\mathbf{p})}\cdot(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}}{\left\|(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|^{2}}=\frac{(\boldsymbol{\chi}_{\rho})_{\mathcal{I}(\mathbf{p})}\cdot(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}}{\tilde{p}(\mathbf{p})}.$ (3)
@@ III Optimization for general states
Input: POVM $\boldsymbol{\Pi}=\{\Pi_{U,\mathbf{b}}\}_{U,\mathbf{b}}$ and default distribution $p$
@@
$\begin{aligned} \mathop{\text{minimize}}_{\omega,t}\quad&\sum_{U}p(U)t_{U}\\ \text{subject to}\quad&-t_{U}\leq\omega_{U,\mathbf{b}}\leq t_{U},\quad\forall U,\mathbf{b},\\ &\sum_{U,\mathbf{b}}\omega_{U,\mathbf{b}}\Pi_{U,\mathbf{b}}=O.\end{aligned}$
@@
OASIS-GT provides a general solution for random targets. However, as will be demonstrated in Section V, applying OASIS-GT to highly structured states such as the GHZ and W states shows diminished performance when compared with previous algorithms such as G-DFE and C-DFE. This is because for these special states, the Pauli operators with nonzero probabilities are sparse and can be grouped efficiently, which eliminates the potential advantage of optimizations that struggle to recognize this property (for a more detailed discussion, see Remark 1 in Appendix B). In this regard, we remark that OASIS-GT is not universally preferable. While it achieves the best performance for Haar-random states, which exhibit minimal structure, we leave a complete characterization of the regimes in which it is most effective to future work.
ALGORITHM 2: OASIS-GHZ.
@@
     $U\leftarrow\bigotimes_{i=1}^{n}U_{\mathbf{p}_{i}}$; $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
@@
     $U\leftarrow\bigotimes_{i=1}^{n}U_{\mathbf{p}_{i}}$; $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
@@ IV Optimization for structured states
In this section, we present efficient and scalable algorithms for structured states, specifically the GHZ and W states. Heuristic algorithms such as G-DFE show scalability issues due to the exponential resource requirements in the grouping procedure and the memory needed to store the groups. To overcome these limitations, we propose more efficient estimators by leveraging non-heuristic grouping strategies. While this work focuses on algorithmic instantiations for GHZ and W states, the underlying optimization framework is general in spirit. Specifically, it relies on the existence of a compact description of the Pauli operators that carry nonzero weight, together with an explicit partition into QWC groups. We expect that analogous algorithms can be developed for other classes of structured states, albeit with additional problem-specific design efforts.
IV.1 OASIS-GHZ
@@ IV.2 OASIS-W
IV.3 Improvement for the W state
Fewer groups generally lead to smaller mean squared error (MSE), since more Pauli observables can be estimated from each measurement shot. In this sense, Algorithms 2 and 3 have the desirable property of minimizing the number of groups.
This grouping is valid, since all of these strings commute with $YYYZ$, but it is suboptimal and redundant because $YYYZ$ itself has zero probability. Consequently, this grouping does not minimize the number of groups, as shown in Table 5 in Appendix F. On the other hand, OASIS-W (and OASIS-GHZ) minimizes the number of groups while ensuring that each group contains no redundant Pauli strings, which accounts for the improvement achieved by our method.
@@ V.1 Setting
We conducted numerical experiments on Haar-random, GHZ, and W states for systems of 3, 4, 5, and 6 qubits to demonstrate the performance of the proposed algorithms. For all targets, we evaluated the proposed OASIS-GT method. In addition, we applied OASIS-GHZ to the GHZ states and OASIS-W to the W states. In each experiment, the target density matrix is denoted by $O$, and the unknown state is modeled as a depolarized version
@@ V.2 MSE comparison
Table 1: Comparison of MSE (1e-4) for Haar-random states. C-DFE and OASIS-GHZ/W are inapplicable, as they are designed for structured targets.
@@
OASIS-W 2.77 2.35 1.40 .721
In Tables 2 and 3, the notion of “scalability” refers to the asymptotic scaling of the overall algorithmic cost with respect to the number of qubits $n$. In particular, the symbols and indicate whether the corresponding method admits polynomial-time or exponential-time scaling in $n$, respectively, taking into account both measurement and classical processing overhead (the polynomial measurement overhead for GHZ/W states follows from the bounds derived in [16]). For the GHZ case in Table 2, although G-DFE and OASIS-GHZ achieve the same performance due to identical grouping structures, they differ substantially in classical computational cost. Specifically, G-DFE is marked as non-scalable, whereas OASIS-GHZ is marked as scalable. Further details on this distinction are provided in Remark 3 in Appendix C.
For OASIS-GT and C-DFE, the user can directly specify the number of measurement shots, and in all reported settings this number is matched to that of G-DFE (see Table 4 in Appendix F) for a fair comparison. In contrast, for OASIS-GHZ/W, the user specifies $\epsilon$ and $\delta$; the total number of shots then becomes a random variable, and is comparable on average to that of G-DFE. Since the MSE typically scales inversely with the number of shots, we apply a correction factor so that the reported MSEs for OASIS-GHZ/W reflect the same average number of shots as G-DFE.
Optimization for Haar-random states was performed using an IOC-POVM based on uniform Pauli measurements, ensuring a consistent basis with G-DFE, which also employs Pauli measurements. The results show that for Haar-random states, OASIS-GT consistently outperforms G-DFE in terms of MSE. For the GHZ states, OASIS-GHZ achieves the best performance for systems of three qubits or larger, while for the W states, OASIS-W achieves the best performance for systems of four qubits or larger. More importantly, the key advantage of OASIS-GHZ/W lies in its scalability. That is, it provides a grouping-based estimator that remains practical for large systems.
@@ VI Discussion
We have introduced a framework for DFE based on operator-aware importance sampling. By formulating estimator optimization as a linear program over an IOC-POVM, OASIS-GT extends and improves upon existing DFE approaches. Furthermore, for structured targets such as the GHZ and W states, the proposed OASIS-GHZ and OASIS-W provide improved fidelity estimation without exponential storage overhead.
A limitation of OASIS-GT is that, like G-DFE, it does not scale to many qubits for Haar-random states because most random states lack the structure necessary for an optimal DFE protocol to be formulated and implemented efficiently. Although LPs are polynomial-time solvable in problem size, the exponential growth of the problem dimension makes the overall computation exponential in $n$. Nevertheless, OASIS-GT provides a principled foundation for operator-aware importance sampling, and enhancing the scalability of the underlying optimization remains an important challenge.
While OASIS-GHZ/W demonstrate both strong performance and scalability, it still inherits a limitation in that it is currently applicable only to Pauli measurement settings. Extending the framework to accommodate more general measurement families remains an important direction for future work.
Additional promising directions include enhancing OASIS-GT through improved surrogate formulations and by considering other IOC-POVMs, and developing target-specific algorithms beyond the GHZ and W states.
Acknowledgements.
This work is in part supported by the National Research Foundation of Korea (NRF, RS-2024-00451435 (20%), RS-2024-00413957 (20%)), Institute of Information & communications Technology Planning & Evaluation (IITP, RS-2021-II212068 (10%), RS-2025-02305453 (15%), RS-2025-02273157 (15%), RS-2025-25442149 (10%), RS-2021-II211343(10%)), a grant funded by the Ministry of Science and ICT (MSIT), Institute of New Media and Communications (INMAC), and the BK21 FOUR program of the Education, Artificial Intelligence Graduate School Program (Seoul National University), and Research Program for Future ICT Pioneers, Seoul National University in 2025.
Data availability
The data that support the findings of this article are openly available [8].
@@ Appendix B Derivation of Algorithm 1
Let $p(U)$ denote the default distribution of $U$ that defines $\boldsymbol{\Pi}$. In other words, we apply $\rho\mapsto U\rho U^{\dagger}$ with probability $p(U)$. Then the POVM elements are
@@
Let us denote the target state’s density matrix as $O$. Then our goal can be formulated as estimating $\text{tr}(\rho O)$ given many copies of an unknown state $\rho$. Since $\boldsymbol{\Pi}$ is an IOC-POVM, $O$ can be expressed as a linear combination of the elements of $\boldsymbol{\Pi}$ as
@@
$\mathop{\text{minimize}}_{\omega}$ $\quad\sum_{U}p(U)\max_{\mathbf{b}}|\omega_{U,\mathbf{b}}|$ (5)
@@
$\mathop{\text{minimize}}_{\omega,t}$ $\quad\sum_{U}p(U)t_{U}$ (6)
@@
Remark 1.
For a Haar-random state, with probability one the Pauli expansion is essentially dense, i.e., almost all Pauli coefficients are nonzero. As a result, both G-DFE and OASIS-GT effectively must sample from (nearly) the full set of $3^{n}$ local Pauli measurement settings. In this regime, OASIS-GT can outperform by globally optimizing its estimator to reduce a variance upper bound rather than relying on G-DFE’s greedy grouping. In contrast, for highly structured targets such as GHZ and W, most Pauli coefficients are exactly zero and the nonzero support is concentrated in a small subset. G-DFE directly exploits this sparsity by focusing only on the nonzero-probability terms and grouping them efficiently. OASIS-GT, however, optimizes a worst-case variance bound over an overcomplete POVM expansion, which need not assign zero weight to many irrelevant Pauli strings. Consequently, it may continue to spread effort over measurement settings that are redundant for these structured targets, and can therefore fail to outperform (or even underperform) G-DFE in this setting.
Remark 2.
Both G-DFE and OASIS-GT incur $2^{\Theta(n)}$ offline (constructing the fidelity estimator) cost for general target states. However, we highlight that the practical utility of our approach lies in the online (the actual fidelity estimation) phase of repeated experimental implementations. In G-DFE, each shot samples a group and produces a scalar contribution that involves computing a weighted sum over the eigenvalues associated with the sampled group [1]. The paper does not explicitly isolate the per-shot classical time of this computation, but it is clear that the online classical processing entails a nontrivial group-dependent weighted sum (whose evaluation scales with the group size). In contrast, OASIS-GT returns the estimator in its most direct and deployment-friendly form. Once the optimization produces $q(U)$ and the estimator $S(U,\mathbf{b})$, the online procedure can be implemented with an alias table [32, 33, 31] for sampling $U\sim q(U)$ and a lookup table for $S(U,\mathbf{b})$. Consequently, the online classical compute per shot is essentially $O(1)$. This difference is particularly beneficial in the common regime where the same optimized fidelity estimator is reused arbitrarily many times. In such cases, the offline cost is amortized, and one repeatedly gains not only the primary benefit of reduced variance in the online fidelity estimate, but also a potentially lower per-measurement classical processing overhead.
@@ Appendix C Derivation of Algorithm 2
$m_{\mathbf{p}}=\left\lceil\frac{2\left\|(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|_{1}^{2}}{\left\|(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right\|^{4}dl\epsilon^{2}}\ln\frac{2}{\delta}\right\rceil$ (11)
@@
In branch 2, $(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\in\mathbb{R}^{1}$ and $\left|(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}\right|=1/\sqrt{d}$, so
@@ Appendix E Derivation of Algorithm 3
In branch 2, $|\mathcal{I}(\mathbf{p})|=d/4$ and all entries of $(\boldsymbol{\chi}_{O})_{\mathcal{I}(\mathbf{p})}$ are $2/n\sqrt{d}$. Therefore,
@@ Appendix F Additional tables
The following tables provide additional numerical data for our experiments.
@@
Table 5: Number of groups produced by G-DFE, OASIS-GHZ, and OASIS-W.